The general solution to equation y - 2y - 3y=0 is a. y=1e3! + ce- b. y=ce + ce-1 C. y = c + c2e- d. none of the above
Answer #1

Note: ) Tx consider and onder linear differential equn with constant coefficients as, a, a, dy ay +9y+Qzy=0 Auxiliary equnsoluna differential equn Y-24-3y=0 (0) Auxiliary equr of (1) is m²_am-3=0 m-3m tm-3=0 mm-3)+1( m-3)=0 =) (M-3) (m+1)=0 ㅋ m=